HIGH–RESOLUTION FOURIER TRANSFORM INFRARED SPECTROSCOPY OF SMALL BORON– CONTAINING MOLECULES

\underline{G. LI} and P. F. BERNATH, Department of Chemistry, University of York, Heslington, York YO10 5DD.

A series of small boron-containing molecules were synthesized in the gas phase using a tube furnace. High-resolution spectra of these species were recorded in either emission or absorption in the mid-infrared region using a Bruker IFS-125HR spectrometer. Our observations contain vibration-rotation bands of BO, the v_{1} and v_{3} bands of HBO, the v_{1} and v_{3} bands of HBS, the v_{1} band of FBO, and the v_{1} band of HBF_{2}. The vibrational bands of HOBO, BF_{2}OH and other boron-containing molecules may also be present. Ab initio calculations were performed at the MRCI level to assist in the vibrational assignments. Preliminary assignments of the spectra for these species will be reported.